Using molarity to calculate grams of solute. What is the concentration of ammonia in a solution made by dissolving 3.75 g of ammonia in 120.0 L of water?
Answer #1

mol = mass/m.wt = 3.75/17.03 = 0.22 mol
Molarity = mol / Volume(L) = 0.22/120.0 = 1.84 * 10^-3
c)1.84